If a current of 0.25 Amperes is applied to a metal object for 30 minutes in a solution of silver nitrate, it will deposit 0.503 grams of silver metal. If the same current is applied for the same amount of time in a solution of nickel sulfate, how much nickel will be deposited? Show your work
Answer #1

The amount of charge = current x time in seconds

= 0.25 x 30 x 60

= 450 Coloumbs

96487 colombs are required for 1 mol of electrons

So 450 colombs = 450 / 96487

= 0.00466 mol of electrons

Now nickel sulfate is NiSO4

Each mol of Ni2+ will require 2 mol of electrons

So here the number of mol of Nickel to get deposited = 0.00466 / 2 = 0.00233 mol

Mass of nickel deposited = number of mole of Nickel x molar mass

= 0.00233 x 58.6934

= 0.137 grams